How many grams of O2 (g) are needed to completely burn 63.1 g C3H8?

C3H8 + 5O2 = 3CO2 + 4H2O

I got 228.986 and I keep getting the wrong answer. What am I doing wrong?
1
Expert's answer
2019-04-03T05:02:11-0400

n= m/Mr = 63.1 / 44 = 1.434 mol (C3H8) and O2 n= 1.434 x 5 = 7.17 mol

m= n x Mr= 7.17 x 32 = 229.454546 g (O2)